The dot products of vector with the vectors (i+j-3k), (i+3j-2k) and (2i+j+4k) are 0,5,8 respectively. Find the vector.
 
 

Grade:12

1 Answers

Shikhar
6 years ago
Let the vector be a ^i + b ^j + c ^kSince dot product of two vectors written in analytical form is a .a`+ b.b` + c.c` Thus , We will obtain three equationa+b-3c=0a+3b-2c=52a +b+4c=8On soving,a=1,b=2,c=1The vector is ^i +2^j+c^k